OpenStudy (anonymous):

What is the sum of a 10-term arithmetic sequence where the first term is 3 and the last term is 75?

OpenStudy (anonymous):

390

OpenStudy (angela210793):

y10=y1+(10-1)*d 75=3+9d 9d=72 d=8 Sn=(y1+y10)*n/2=(3+75)*10/2=78*5=390

OpenStudy (anonymous):

formula used n(a+l)/2 n=the number of terms a=1st term l=last term

OpenStudy (anonymous):

sritama was correct, thank you :):)